40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 721     Entry time: Wed Jul 23 10:49:37 2008
Author: Max 
Type: Update 
Category: Computer Scripts / Programs 
Subject: Weekly Progress Report 
This week I installed the magnetometer. The channels seem to be reading correctly. I'm back to working on noise budget and have added the MICH and will soon add the PRC source. The various source-specific scripts still need to be adjusted and the transfer functions remeasured since they do not match in any reasonable manner the SRD Rana put out in the e-log yesterday.
ELOG V3.1.3-